# From Birdwatch to Community Notes, from Twitter to X: four years of community-based content moderation

### v2

@@ INTRODUCTION
The rapid production and spread of user-generated content on social media platforms in the absence of editorial oversight has increased users’ exposure to false or misleading information1. In response, social media platforms have taken various approaches to content moderation. Content Moderation refers to the process of monitoring, flagging, or removing content that violates community guidelines or is deemed harmful. The main moderation strategies have been expert review and algorithmic systems. However, each comes with limitations. Expert evaluation, while often accurate, is costly and impractical at scale, given the sheer volume of content that needs moderation2. Automated algorithmic methods, on the other hand, are constrained by the biases and quality of their training data, often reinforcing systemic biases in classification3. In light of these challenges, community-based content moderation, leveraging the collective judgment of users, has emerged as a promising alternative.
On January 23, 2021, X (formerly Twitter) launched Community Notes (formerly Birdwatch), the first large-scale community-driven initiative to moderate misleading content. To participate, Contributors must have been active on X for at least six months, have a verified phone number, and maintain a clean record of rule compliance4.
Contributors can write a Note to provide the missing context in relation to the Posts they find misleading. Other Contributors then rate these Notes as helpful, somewhat helpful, or unhelpful (Figure 1 (a)). The Note that receives the Helpful status based on the ratings from a diverse group of Contributors is displayed beneath the original Post in the timeline (an example is shown in Figure 1 (b))5.
Each Contributor has two impact metrics: writing impact and rating impact. The former increases when a Contributor’s Notes are consistently rated as helpful. The latter increases when the Contributor’s rating of a Note aligns with the eventual status of the Note based on the rating algorithm, and decreases otherwise6. New Contributors begin by rating existing Notes and can only write their own Notes once their rating score reaches a threshold of 57. Afterwards, the Contributors can decide to write a Note on any Post. Furthermore, users on X can request a Note for a Post; if a Post receives enough requests, top Contributors, those with high writing impact, are notified to draft a Note for that Post8.
@@
Community Notes has experimented with various rating systems and decision rules to determine which note should be displayed for each post. The description provided above outlines only the core functionality of the system as of the time of writing; we do not aim to detail the specific algorithms and configurations, as these continue to evolve. In the pilot version of Community Notes, Notes that received the highest number of helpful ratings were picked to be shown in the timeline, regardless of the raters’ past behaviour. This was not immune to polarised rating among Contributors9 and partisan behaviour, such as Contributors rating Notes written by co-partisans as helpful and labelling Posts from cross-partisan Contributors as misleading10. To mitigate this issue, Community Notes adopted a bridging algorithm11 in the rating system. This approach places Contributors along an opinion spectrum based on their historical rating behaviour. A Note is deemed helpful only if it receives enough positive ratings from Contributors with opposing viewpoints. The goal of this algorithm is to elevate Notes that are rated as helpful by a diverse set of Contributors. Community Notes is entirely open-source, with both the code and data available to the public12.
@@
Several studies have assessed the effectiveness of Community Notes in comparison with other content moderation methods. Findings indicate that it is effective at identifying misinformation, frequently flagging misleading content before expert fact-checkers. There is substantial agreement between Community Notes Contributors and professional fact-checkers, with their classification aligning closely with expert evaluations13. Compared to expert-led moderation, Community Notes is also more scalable and cost-effective14. Unlike other crowdsourced approaches, such as “snoping”, in which users respond directly to Posts with fact-checks, Community Notes Contributors tend to prioritise high-visibility content, particularly posts from influential users. When both systems evaluate the same content, a relatively rare occurrence, their assessments typically align15.
@@
Despite these advantages, Community Notes has notable limitations. It often operates more slowly than approaches such as “snoping”15. A recent study analysing over 2.2 million Posts found that 99.3% of misleading Posts received debunking comments within two hours of publication, whereas it takes an average of 24.29 hours for a Community Note to appear beneath a Post16. Moreover, Community Notes’ open design leaves it vulnerable to manipulation and adversarial attacks13. It also relies heavily on the work of professional fact-checkers: one in twenty Notes explicitly references fact-checking sources, with this proportion increasing for sensitive topics17. These findings indicate that the production of high-quality Notes remains closely dependent on the broader fact-checking ecosystem.
@@
While the studies above demonstrate that Community Notes can accurately identify misleading content and produce high-quality Notes, a critical question remains: can Community Notes reduce user engagement with such content? Results from A/B testing indicate that users exposed to Community Notes are 25–34% less likely to like or share flagged Posts11. Additional studies report that Posts with attached Notes have higher deletion rates18, and that attached Notes reduce reposts. This effect is more pronounced for Posts containing embedded media than for text-only Posts19. In addition to reposts, an attached Note reduces the number of likes, views, and replies of a Post. Furthermore, when Notes appear beneath a Post, that Post tends to spread less widely and deeply across the platform20.
@@
Community Notes also shape how readers respond to misinformation. Kankham and Hou (2024) find that Notes are particularly effective in reducing belief and the spread of “wish” rumours (misinformation aligned with users’ desires), whereas presenting related news articles is more effective in countering “dread” rumours (misinformation that evokes fear)21. Users generally view Community Notes more favourably than other interventions, such as misinformation flags. Unlike binary labels that simply mark content as false, Notes provide explanatory context, which fosters greater trust and acceptance. Drolsbach et al. (2024) show that this contextual approach is more persuasive, as users tend to prefer interpretive frameworks over authoritative declarations of truth22.
@@
However, recent studies indicate that the overall rollout of Community Notes has not substantially curtailed engagement with misleading content. This limitation is attributed mainly to the delay in Notes reaching “helpful” status, often well after a Post’s most viral phase23, 24. Renault et al. (2024) report that while annotated Posts experience nearly a 50% drop in reposts and over a 30% reduction in replies and quotes, this effect is highly temporal. On average, it takes 15 hours for a Note to be published, by which time a Post has typically reached 80% of its total audience18. Similarly, Chuai et al. (2024) report an even longer delay—an average of 75.5 hours between post creation and note attachment—by which time 96.7% of reposts have already occurred 19. De et al. (2024) report that in 91% of Posts where at least one Note was proposed, none reached “helpful” status25. This delay is particularly problematic in time-sensitive or politically charged contexts. For instance, one report found that 74% of accurate Notes related to the 2024 U.S. presidential election were never shown to users, allowing misleading Posts without Notes to spread 13 times faster than those annotated with Community Notes26. Researchers attribute this to the design of the rating algorithm: the more polarising the content (e.g., national elections), the less likely accurate Notes are to receive “helpful” status27. Related to this is the report that suggests most contributors have yet to produce a single Note rated as helpful28.
@@
Community Notes influence both posting behaviour and the ways users engage with content on the platform. Studies find that users whose content receives Notes tend to Post less overall, yet their subsequent Posts exhibit greater cognitive processing. Additionally, exposure to Community Notes improves reliability in users’ own writing, while reducing extreme sentiment. However, these cognitive benefits are accompanied by decreased overall activity, as users exposed to Community Notes participate less frequently29. Receiving a Note can also have the unintended effect of increasing a user’s visibility: users, particularly those with smaller followings, who are “noted” often gain additional followers, potentially amplifying their reach rather than diminishing it28.
@@
Several studies have examined the characteristics of Notes produced in Community Notes to understand better what makes a Note more effective or persuasive. Research indicates that Notes on Posts that are flagged as misleading tend to be longer, more complex, and more negatively worded; traits that may make them more informative but potentially less accessible to general audiences30. The emotional tone of a Note also plays a key role: while emotionally charged language can increase engagement, it tends to reduce perceived trustworthiness. On average, Notes with high emotional content are rated as less helpful than those adopting a more neutral tone31. Moreover, Notes that closely align with the topic of the associated Post receive higher helpfulness ratings32.
@@
Citing sources is another critical element of Note quality. Contributors are encouraged to support their claims with external references, and including a source increases the likelihood of a Note reaching “helpful” status by a factor of 2.33. However, sources perceived as politically biased reduce the likelihood of a Note being rated as helpful33. Kangur et al. (2024) analysed citation patterns and found that the most frequently referenced sources were X and Wikipedia, with a noticeable left-leaning bias. Notes that cited more balanced or factually rigorous sources received higher helpfulness scores, reaffirming the other findings34.
@@
Contributors select which Posts to write Community Notes for based on multiple factors. A recent study shows that among 90,000 Posts for which X users requested a Note, the Posts perceived as more misleading by GPT-4.1 and the Posts by authors who are more frequently fact-checked, were more likely to attract Notes35. Political partisanship also plays a central role in how Contributors both select and evaluate Notes. Prior research shows that partisanship is often a stronger predictor of judgment than the content itself. Contributors are more likely to evaluate Posts from political opponents negatively and to rate Notes written by Contributors of opposing affiliations as unhelpful, thereby reinforcing existing ideological divides10. Network analyses of positive interactions reveal that the initial rating system, where the Note with the highest number of helpful ratings was surfaced, did not promote cross-partisan support. Instead, Contributors clustered into polarised groups9. This pattern mirrors broader trends of political polarisation observed on social media. Studies of signed interaction networks in Community Notes reinforce this finding, showing that Contributors consistently form clusters based on shared political ideologies36. Topic modelling and network analyses further demonstrate that such polarisation is especially pronounced in political discussions, whereas evaluations of non-political Posts display substantially less ideological bias37.
@@
Despite this, there is evidence that diversity within the pool of fact-checkers can improve fact-checking outcomes. Research comparing individual tagging (e.g., labelling misinformation independently) with Community Notes finds that the latter exposes users to more diverse perspectives. Unlike individual tagging, which can deepen echo chambers, Community Notes’ peer-review structure temporarily increases informational diversity38. Although this effect is not lasting, it underscores the potential of collective moderation to mitigate polarisation when thoughtfully implemented. Research suggests that Community Notes can improve the quality of public discourse and support misinformation detection, drawing parallels to deliberative platforms such as Polis, which structure participation to promote reflection and consensus39.
@@
The system’s evolution has also influenced Contributors’ behaviour over time. Early in the program, most Notes were written on Posts from Contributors estimated to lean liberal28. However, as the Contributor base broadened, the partisan distribution shifted. A 2025 study by Renault et al. found that Posts from Republican users were more likely to be flagged with Community Notes, reflecting ongoing changes in participation and focus40.
@@
While Community Notes offer scalable fact-checking and collective decision-making benefits, it also carries risks, including system manipulation, exploitation of unpaid labour, and the marginalisation of underrepresented voices41. Furthermore, some features—such as prompting Contributors to revisit flagged content—may unintentionally reinforce exposure to misinformation42. Other ethical concerns have also been raised about the open structure of Community Notes, particularly the potential psychological harm to Contributors who encounter harmful or sensitive content43.
Several studies have proposed modifications to Community Notes to enhance its effectiveness. Lloyd et al. (2025)44 argue that researchers should focus on how the design and implementation of such systems influence their efficiency, including features such as interaction between Contributors, Note presentation, and the integration of technological advances such as Large Language Models (LLMs). Various approaches have already been explored. HawkEye, for instance, is a graph-based algorithm that jointly evaluates the quality of Contributors, Notes, and Posts through repeated iterations. It applies basic scoring principles and a smoothing technique to address cases with limited data, thereby improving the reliability and consistency of its evaluations45. Supernotes adopt a different approach by leveraging large language models (LLMs) to generate Notes that are more likely to be rated as “helpful” than those written by individual Contributors. In this framework, an LLM reads all existing Notes, generates multiple candidate Notes, and evaluates them with a Personalised Helpful Model (PHM), which estimates the probability of receiving a “helpful” rating from a synthetic jury of diverse raters. The highest-scoring Note is then presented as the Supernote, and empirical evidence shows that Contributors rate these as significantly more helpful than the best human-written Notes25. Some scholars extend this idea into a human–AI hybrid framework in which both human Contributors and LLMs can propose Notes, but only humans serve as raters and evaluators. This division of labour preserves the legitimacy of human judgment while enabling LLMs to accelerate the generation of high-quality Notes46. Pushing this idea further, Costabile et al. (2025) simulated crowds using generative agents with diverse demographics and ideological perspectives. These agents outperformed human crowds in truthfulness classification, demonstrated higher internal consistency, were less susceptible to social and cognitive biases, and relied more on systematically informative criteria47.
@@
A second line of research emphasises diversity and collaboration in the Note-writing process rather than focusing solely on the rating stage9. In an online experiment, Juncosa et al. found that duos produced more helpful Notes than individuals, with diverse duos performing particularly well on Republican-leaning Posts48. Building on this insight, Mohammadi and Yasseri (2025) used LLMs to synthesise political diversity during Note writing. In a study involving over 800 participants, Contributors who received argumentative feedback from an LLM were more likely to produce higher-quality Notes49.
@@
In this work, we collected all Community Notes and corresponding ratings published over a four-year period, from January 23, 2021, to January 23, 2025. We applied language detection to filter for English-language Notes and, on this subset, conducted topic modelling, URL and domain extraction, and constructed interaction networks between Contributors based on their ratings of each other’s Notes. All processed data—including annotated Notes, interaction networks, and analysis outputs—are publicly released to support future research (see *Resource Availability* section). Although our present analysis is limited to English, the accompanying codebase is readily adaptable for multilingual extensions.

### v3

@@ INTRODUCTION
The rapid production and spread of user-generated content on social media platforms in the absence of editorial oversight has increased users’ exposure to false or misleading information (Thai et al. (2016)). In response, social media platforms have taken various approaches to content moderation. Content Moderation refers to the process of monitoring, flagging, or removing content that violates community guidelines or is deemed harmful. The main moderation strategies have been expert review and algorithmic systems. However, each comes with limitations. Expert evaluation, while often accurate, is costly and impractical at scale, given the sheer volume of content that needs moderation (Hassan et al. (2015)). Automated algorithmic methods, on the other hand, are constrained by the biases and quality of their training data, often reinforcing systemic biases in classification (Binns et al. (2017)). In light of these challenges, community-based content moderation, leveraging the collective judgment of users, has emerged as a promising alternative.
On January 23, 2021, X (formerly Twitter) launched Community Notes (formerly Birdwatch), the first large-scale community-driven initiative to moderate misleading content. To participate, Contributors must have been active on X for at least six months, have a verified phone number, and maintain a clean record of rule compliance (X Documentation (2025g)).
Contributors can write a Note to provide the missing context in relation to the Posts they find misleading. Other Contributors then rate these Notes as helpful, somewhat helpful, or unhelpful (Figure 1(a)). The Note that receives the Helpful status based on the ratings from a diverse group of Contributors is displayed beneath the original Post in the timeline (an example is shown in Figure 1(b)) (X Documentation (2025d)).
Each Contributor has two impact metrics: writing impact and rating impact. The former increases when a Contributor’s Notes are consistently rated as helpful. The latter increases when the Contributor’s rating of a Note aligns with the eventual status of the Note based on the rating algorithm, and decreases otherwise (X Documentation (2025e)). New Contributors begin by rating existing Notes and can only write their own Notes once their rating score reaches a threshold of 5 (X Documentation (2025b)). Afterwards, the Contributors can decide to write a Note on any Post. Furthermore, users on X can request a Note for a Post; if a Post receives enough requests, top Contributors, those with high writing impact, are notified to draft a Note for that Post (X Documentation (2025f)).
@@
Community Notes has experimented with various rating systems and decision rules to determine which note should be displayed for each post. The description provided above outlines only the core functionality of the system as of the time of writing; we do not aim to detail the specific algorithms and configurations, as these continue to evolve. In the pilot version of Community Notes, Notes that received the highest number of helpful ratings were picked to be shown in the timeline, regardless of the raters’ past behaviour. This was not immune to polarised rating among Contributors (Yasseri and Menczer (2023)) and partisan behaviour, such as Contributors rating Notes written by co-partisans as helpful and labelling Posts from cross-partisan Contributors as misleading (Allen et al. (2022)). To mitigate this issue, Community Notes adopted a bridging algorithm (Wojcik et al. (2022)) in the rating system. This approach places Contributors along an opinion spectrum based on their historical rating behaviour. A Note is deemed helpful only if it receives enough positive ratings from Contributors with opposing viewpoints. The goal of this algorithm is to elevate Notes that are rated as helpful by a diverse set of Contributors. Community Notes is entirely open-source, with both the code and data available to the public (X Documentation (2025c)).
@@
Several studies have assessed the effectiveness of Community Notes in comparison with other content moderation methods. Findings indicate that it is effective at identifying misinformation, frequently flagging misleading content before expert fact-checkers. There is substantial agreement between Community Notes Contributors and professional fact-checkers, with their classification aligning closely with expert evaluations (Saeed et al. (2022)). Compared to expert-led moderation, Community Notes is also more scalable and cost-effective (Martel et al. (2024)). Contributors tend to focus on posts from influential users with larger followings, whereas experts typically target content from less influential accounts (Drolsbach and Pröllochs (2023b)).Unlike other crowdsourced approaches, such as “Snoping”, in which users respond directly to Posts with fact-checks, Community Notes Contributors tend to prioritise high-visibility content, particularly posts from influential users. When both systems evaluate the same content, a rare occurrence, their assessments typically align (Pilarski et al. (2024)).
@@
Despite these advantages, Community Notes has notable limitations. It often operates more slowly than approaches such as “snoping” (Pilarski et al. (2024)). A recent study analysing over 2.2 million Posts found that 99.3% of misleading Posts received debunking comments within two hours of publication, whereas it takes an average of 24.29 hours for a Community Note to appear beneath a Post (Zhang et al. (2025)). Moreover, Community Notes’ open design leaves it vulnerable to manipulation and adversarial attacks (Saeed et al. (2022)). It also relies heavily on the work of professional fact-checkers: one in twenty Notes explicitly references fact-checking sources, with this proportion increasing for sensitive topics (Borenstein et al. (2025)). These findings indicate that the production of high-quality Notes remains closely dependent on the broader fact-checking ecosystem.
@@
While the studies above demonstrate that Community Notes can accurately identify misleading content and produce high-quality Notes, a critical question remains: can Community Notes reduce user engagement with such content? Results from A/B testing indicate that users exposed to Community Notes are 25–34% less likely to like or share flagged Posts (Wojcik et al. (2022)). Additional studies report that Posts with attached Notes have higher deletion rates, and that attached Notes reduce reposts. This effect is more pronounced for Posts containing embedded media than for text-only Posts (Chuai et al. (2024)). Similarly, Drolsbach and Pröllochs (2023b) estimate that misleading Posts with attached Notes receive 36.85% fewer reposts than non-misleading Posts. Beyond reposts, annotated Posts tend to accumulate fewer likes, views, and replies overall. Moreover, when Notes appear beneath a Post, that Post spreads less widely and deeply across the platform (Slaughter et al. (2025)). However, being fact-checked through Community Notes does not appear to meaningfully reduce an author’s follower count, suggesting that the system’s influence on social connections remains limited (Bobek and Pröllochs (2025)).
@@
Community Notes also shape how readers respond to misinformation. Studies show that once a Note is displayed underneath Post, there is a measurable shift in users’ emotional responses, including a 7.3% increase in negativity in replies. More specifically, displayed Notes are cause a 13.2% increase in anger, a 4.7% increase in disgust, and a 16.0% rise in moral outrage in replies. These effects are more pronounced for political content than for non-political content (Chuai et al. (2025a)). Kankham and Hou (2024) find that Notes are particularly effective in reducing belief and the spread of “wish” rumours (misinformation aligned with users’ desires), whereas presenting related news articles is more effective in countering “dread” rumours (misinformation that evokes fear). Users generally view Community Notes more favourably than other interventions, such as misinformation flags. Unlike binary labels that simply mark content as false, Notes provide explanatory context, which fosters greater trust and acceptance. Drolsbach et al. (2024) show that this contextual approach is more persuasive, as users tend to prefer interpretive frameworks over authoritative declarations of truth.
@@
However, recent studies indicate that the overall rollout of Community Notes has not substantially curtailed engagement with misleading content. This limitation is attributed mainly to the delay in Notes reaching “helpful” status, often well after a Post’s most viral phase (Chuai et al. (2023); Bak-Coleman (2023)). Renault et al. (2024) report that while annotated Posts experience nearly a 50% drop in reposts and over a 30% reduction in replies and quotes, this effect is highly temporal. On average, it takes 15 hours for a Note to be published, by which time a Post has typically reached 80% of its total audience. Similarly, Chuai et al. (2024) show that by the time a Post reaches its half-life, only 13.5% of helpful Notes have been attached. De et al. (2024) further find that in 91% of Posts where at least one Note was proposed, none ultimately reached “helpful” status. Researchers suggest that this pattern is driven in part by insufficient diversity of opinions among raters. Arjmandi-Lari et al. (2025) show that having a Note published increases the retention of first-time Contributors by 5%, implying that low publication rates may threaten the long-term sustainability of the Community Notes ecosystem.
These delays are particularly problematic in time-sensitive or politically charged contexts. For instance, one report found that 74% of accurate Notes related to the 2024 U.S. presidential election were never shown to users, allowing misleading Posts without Notes to spread 13 times faster than those annotated with Community Notes (Center for Countering Digital Hate (2024)). Researchers attribute this to the design of the rating algorithm: the more polarising the content (e.g., national elections), the less likely accurate Notes are to receive “helpful” status (Bouchaud and Ramaciotti (2025)). Related to this is the report that suggests most contributors have yet to produce a single Note rated as helpful (Wirtschafter and Majumder (2023)).
The system exhibits distinctive patterns in cases where Contributors write Notes asserting that a Post should not be considered misleading. These Notes often begin with “NNN,” meaning “Note Not Needed”. Razuvayevskaya et al. (2025) argue that such cases reflect a repurposing of the platform for debate rather than moderation. Their study shows that Posts receiving an “NNN” Note are more likely to yield a published Note than Posts that do not, suggesting that Contributors use the system not only to correct misleading content but also to contest whether moderation is warranted.
@@
Community Notes influence not only how users engage with content but also how they post on the platform. Studies show that users whose content receives Notes tend to post less frequently overall; however, their subsequent Posts display greater cognitive processing. Exposure to Community Notes also increases the reliability of users’ own writing while reducing extreme sentiment. These cognitive benefits, however, come with a trade-off as users exposed to Notes tend to participate less actively on the platform (Borwankar et al. (2022)). Receiving a Note can also have the unintended effect of increasing a user’s visibility: users, particularly those with smaller followings, who are “noted” often gain additional followers, potentially amplifying their reach rather than diminishing it (Wirtschafter and Majumder (2023)).
At the same time, research examining how Community Notes Contributors create content on X finds that system-level changes on Community Notes can shape their behaviour of users on X. After the introduction of a privacy-preserving policy that anonymised Contributor identities—preventing their Community Notes activity from being linked to their X profiles—Contributors began posting more frequently, and the neutrality and overall quality of their Posts improved (Borwankar et al. (2024)).
@@
Several studies have examined the characteristics of Notes produced in Community Notes to better understand what makes a Note more effective or persuasive. Research indicates that Notes on Posts that are flagged as misleading tend to be longer, more complex, and more negatively worded; traits that may make them more informative but potentially less accessible to general audiences (Pröllochs (2022)). The emotional tone of a Note also plays a key role: while emotionally charged language can increase engagement, it tends to reduce perceived trustworthiness. On average, Notes with high emotional content are rated as less helpful than those adopting a more neutral tone (Phillips et al. (2024)). Moreover, Notes that closely align with the topic of the associated Post receive higher helpfulness ratings (Simpson (2022)).
@@
Citing sources is another critical element of Note quality. Contributors are encouraged to support their claims with external references, and including a source increases the likelihood of a Note reaching “helpful” status by a factor of 2.33. However, sources perceived as politically biased reduce the likelihood of a Note being rated as helpful (Solovev and Pröllochs (2025)). Kangur et al. (2024) analysed citation patterns and found that the most frequently referenced sources were X and Wikipedia, with a noticeable left-leaning bias. Notes that cited more balanced or factually rigorous sources received higher helpfulness scores.
@@
Contributors select which Posts to write Community Notes for based on multiple factors. A recent study shows that among 90,000 Posts for which X users requested a Note, the Posts perceived as more misleading by GPT-4.1 and the Posts by authors who are more frequently fact-checked, were more likely to attract Notes (Chuai et al. (2025b)). Political partisanship also plays a central role in how Contributors both select and evaluate Notes. Prior research shows that partisanship is often a stronger predictor of judgment than the content itself. Contributors are more likely to evaluate Posts from political opponents negatively and to rate Notes written by Contributors of opposing affiliations as unhelpful, thereby reinforcing existing ideological divides (Allen et al. (2022)). Network analyses of positive interactions reveal that the initial rating system, where the Note with the highest number of helpful ratings was surfaced, did not promote cross-partisan support. Instead, Contributors clustered into polarised groups (Yasseri and Menczer (2023)). This pattern mirrors broader trends of political polarisation observed on social media. Studies of signed interaction networks in Community Notes reinforce this finding, showing that Contributors consistently form clusters based on shared political ideologies (Fraxanet et al. (2024)). Topic modelling and network analyses further demonstrate that such polarisation is especially pronounced in political discussions, whereas evaluations of non-political Posts display substantially less ideological bias (Champaigne (2022)).
Recent work suggests that the delays in Note publication also stem from such behaviour. Truong et al. (2025b) exposed the algorithm to more than 15,000 simulated datasets with varying attributes and found it to be highly sensitive to rater biases—such as the tendency to downrate Notes written by Contributors of the opposite political persuasion. Their results show that even a small fraction of low-quality raters can suppress helpful Notes to the extent that they never reach publication.
@@
Despite these limitations, evidence suggests that diversity within the pool of fact-checkers can enhance fact-checking outcomes. Studies show that political motivation can help address a key challenge in volunteer-based moderation systems—insufficient content flagging. Contributors with strong political motives flag twice as many Posts as less politically motivated individuals, without any notable decline in the quality of their flags (Martel et al. (2025)). Moreover, research comparing individual tagging (e.g., labelling misinformation independently) with Community Notes finds that the latter exposes users to more diverse perspectives. Unlike individual tagging, which can deepen echo chambers, Community Notes’ peer-review structure temporarily increases informational diversity (Kim et al. (2025)). Although this effect is not lasting, it underscores the potential of collective moderation to mitigate polarisation when thoughtfully implemented. Research suggests that Community Notes can improve the quality of public discourse and support misinformation detection, drawing parallels to deliberative platforms such as Polis, which structure participation to promote reflection and consensus (Megill et al. (2022)).
@@
The system’s evolution has also influenced Contributors’ behaviour over time. Early in the program, most Notes were written on Posts from Contributors estimated to lean liberal (Wirtschafter and Majumder (2023)). However, as the Contributor base broadened, the partisan distribution shifted. Renault et al. (2025) found that Posts from Republican users were more likely to be flagged with Community Notes, reflecting ongoing changes in participation and focus.
@@
While Community Notes offer scalable fact-checking and collective decision-making benefits, it also carries risks, including system manipulation, exploitation of unpaid labour, and the marginalisation of underrepresented voices (Benjamin (2021)). Furthermore, some features—such as prompting Contributors to revisit flagged content—may unintentionally reinforce exposure to misinformation (Wang and Lucas (2024)). Other ethical concerns have also been raised about the open structure of Community Notes, particularly the potential psychological harm to Contributors who encounter harmful or sensitive content (Augenstein et al. (2025)).
Several studies have proposed modifications to Community Notes to enhance its effectiveness. Lloyd et al. (2025) argue that future research should focus on how the system’s design and implementation influence its efficiency, including aspects such as Contributor interaction, Note presentation, and the integration of technological advances like Large Language Models (LLMs). Building on this, various algorithmic and hybrid approaches have been explored.
One such approach is HawkEye, a graph-based algorithm that jointly evaluates the quality of Contributors, Notes, and Posts through iterative computation. By applying scoring principles and smoothing techniques to handle sparse data, HawkEye improves the consistency and robustness of evaluations (Mujumdar and Kumar (2021)).
Supernotes adopt a different approach by leveraging large language models (LLMs) to generate Notes that are more likely to be rated as “helpful” than those written by individual Contributors. In this framework, an LLM reads all existing Notes, generates multiple candidate Notes, and evaluates them with a Personalised Helpful Model (PHM), which estimates the probability of receiving a “helpful” rating from a synthetic jury of diverse raters. The highest-scoring Note is then presented as the Supernote, and empirical evidence shows that Contributors rate these as significantly more helpful than the best human-written Notes (De et al. (2024)).
Additionally, Wu et al. (2025) proposed CrowdNotes+, a similar hybrid framework tailored to health-related content. In this model, LLMs can either autonomously write Notes or use evidence gathered by human Contributors. Results show that LLMs produce more accurate and contextually balanced Notes than humans, and when sourcing their own evidence, they tend to select higher-quality material.
Other scholars extend this idea into a human–AI hybrid framework in which both human Contributors and LLMs can propose Notes, but only humans serve as raters and evaluators. This division of labour preserves the legitimacy of human judgment while enabling LLMs to accelerate the generation of high-quality Notes (Li et al. (2025)). Pushing this idea further, Costabile et al. (2025) simulated crowds using generative agents with diverse demographics and ideological perspectives. These agents outperformed human crowds in truthfulness classification, demonstrated higher internal consistency, were less susceptible to social and cognitive biases, and relied more on systematically informative criteria.
@@
Another line of research emphasises diversity and collaboration in the Note-writing process rather than focusing solely on the rating stage (Yasseri and Menczer (2023)). In an online experiment, Juncosa Calahorrano et al. (2025) found that duos produced more helpful Notes than individuals, with diverse duos performing particularly well on Republican-leaning Posts. Building on this insight, Mohammadi and Yasseri (2025) used LLMs to synthesise political diversity during Note writing. In a study involving over 800 participants, Contributors who received argumentative feedback from an LLM were more likely to produce higher-quality Notes.
Building on these developments, researchers have also begun to use Community Notes as a rich empirical resource for studying broader online behaviours and misinformation dynamics. For example, Drolsbach and Pröllochs (2023a) analysed contributor responses to questions assessing whether misleading posts were perceived as believable and/or harmful. By linking these perceptions to the repost count, they found that misleading posts considered easily believable but not particularly harmful were more likely to be reposted.
@@
In this work, we collected all Community Notes and corresponding ratings published over a four-year period, from January 23, 2021, to January 23, 2025. We applied language detection to filter for English-language Notes and, on this subset, conducted topic modelling, URL and domain extraction, and constructed interaction networks between Contributors based on their ratings of each other’s Notes. All processed data—including annotated Notes, interaction networks, and analysis outputs—are publicly released to support future research (see Data Availability section). Although our present analysis is limited to English, the accompanying codebase is readily adaptable for multilingual extensions.